OpenStudy (s):

What is the inflection point of f(x)=x^2(x^2-4) ?

OpenStudy (anonymous):

\[f(x)=x^2(x^2-4)=x^4-4x^2\]\[f'(x)=4x^3-8x\]\[f''(x)=12x^2-8\]\[12x^2-8=0\implies x=\pm\sqrt{\frac{2}{3}}\]

OpenStudy (anonymous):

Then I can't do much.